The compressed air compressor absorbs air at 101.3 kPa, 10 ° C and relative humidity
60%. After the compressor, the compressed air temperature is 30 ° C and the pressure is 300 kPa. What is the relative humidity of the compressed air produced?
1
Expert's answer
2019-04-08T12:58:01-0400

As far as humidity γ2  = n2/nl2  i.e.  γ2 = p2/pl2 = p1(pt2/pt1)/pl2 = γ1(pl1/pl2) (pt2/pt1) = 0.6(1.23/4.21)(300/101.3) = 0.52 or 52%.